Subject: January 2014 APO 3.5m users committee minutes
From: Suzanne Hawley
Submitted: Mon, 27 Jan 2014 12:05:48 -0800 (PST)
Message number: 1325
(previous: 1324,
next: 1326
up: Index)
APO 3.5-m Users Committee Phonecon, 1/13/14
Attending: Suzanne Hawley, Mark Klaene, Rene Walterbos, Bill Ketzeback,
John Bally, Mary Beth Kaiser, John Wisniewski, Joanne Hughes Clark, Scott
Anderson, Michael Strauss, Rachel Kuzio de Naray, Jon Holtzman, Remy
Indebetouw
Absent: Al Harper, Jeff Bary, Bruce Gillespie, Joe Huehnerhoff, Russet
McMillan
**********************************
User feedback and comments from institutional representatives:
o Washington (Scott Anderson) - Scott mentioned that one user was
concerned with data reduction using the new timestamps. This is an agenda
item discussed later in the meeting.
o Chicago (Al Harper) - no report.
o New Mexico State (Rene Walterbos) - Rene had nothing new to report.
o Princeton (Michael Strauss) - Michael said that everything was fine.
o Colorado (John Bally) - John had nothing new to report.
o Virginia (Remy Indebetouw) - Remy had nothing to report.
o Johns Hopkins (Mary Beth Kaiser) - Mary Beth had no concerns to report.
o Georgia State (Rachel Kuzio de Naray) - Rachel had nothing to report.
o Seattle Univ. (Joanne Hughes Clark) - Joanne said everything was fine.
o U. Oklahoma (John Wisniewski) - No observing yet, nothing to report.
o NAPG (Jeff Bary) - no report.
**********************************
Discussion of telescope/instruments report:
In addition to the normal report, Mark mentioned the 3.5-m policy for snow
and ice removal from the enclosure roof. The engineering staff makes every
effort to remove snow accumulation from the roof as soon as it becomes
safe to do so by shoveling. The exception to this is accumulation of snow
or ice on the four lightning arrestors. These 12 foot poles extend above
the corners of the enclosure roof are difficult and unsafe to get to. Some
preliminary investigation has gone into finding a solution to keep them
free of snow and ice but so far we have been unsuccessful in our search.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope and Instruments Highlights
11/25/13 through 01/07/14
0) Overview
Generally a productive period with a large snowfall of the season now on
the ground that of course came at the start of the NMSU Holiday break.
However with extra effort by the crew, snow removal limited loss time to a
minimum.
The H1N1 flu hit the area hard over the Holiday break affecting day and
night personnel. With support from all the staff we were able to keep the
telescopes open. Most personnel are back up to at least 80%.
Patrick Galume who was filling in for Alaina Bradley while she was on
maternity leave is now hired as a SDSS observer through SDSS III and
Alaina has returned to work.
1) Telescope
Telescope has been working very well. There has been good progress made
on the guider rewrite and some observers have volunteered to test the new
guiding software on sky since engineering time is limited this time of
year. We will continue to make progress on this over the next several
months and when satisfied convert the new guider routines to the default.
2) Instruments
DIS has been operating normally except for a problem with one of the
grating bays. We have isolated the problem and ordered replacement parts
however a moderate amount of dis-assembly will be required to replace the
suspect part.
Agile is operational with no problems
TripleSpec had no operational problems this period.
SPIcam variable dark current has been isolated to the current ion pump
type and controller. We continue to work with the ion pump manufacturer to
come up with a long term solution. Currently we see a spike that may
affect dark current about 1/day and if it occurs during observations we
will let the observer know to check their data. Generally the effect is
small.
Echelle is operational with no problems
NIC-FPS is operating in shared-risk observing mode due to occasional image
corruption issue. NFS less than 9 is operational.
**********************************
o Time stamps in FITS headers:
As of August 30, 2013, the 3.5-m has implemented a TAI time standard in
all the instrument headers. All time keywords are referenced to TAI. In an
effort to make all time keywords the same from instrument to instrument
various keywords are no longer being written to some instrument headers.
The standard keywords are: TIMESYS, UTC-TAI, UT1-TAI, LST, DATE-OBS (TAI
time at start of exposure in year-month-day-hou r:minute:second format),
and EXPTIME. Every effort has been made to give users the minimum number
of time related keywords with which they can recreate any of the old
keywords on their own in post processing.
Some iraf packages have had difficulty with the format of the timestamp
(YYYY-MM-DDTHH:MM:SS.S). Suzanne would be interested to hear if users have
been successful in running their data through their reduction pipelines
since the changes to the fits headers have been made.
==> ACTION ITEM: Committee members please query their users about issues
==> with reducing data due to the new format of the timestamp.
o APO users wiki:
Jon Holtzman is working on setting up an APO 3.5m wiki at NMSU. Still
considering whether access will be via a user list or open to the world.
He gave a description of how such a users wiki for data reduction and
instrument specifics might be used. It was noted that the above issue
about time keywords would be a good item to have on the wiki. Jon will
send an email to users when the wiki is open for business.
o ARCTIC update:
Mark mentioned 2 items out of the PDR. First was a review of the optics
design with some input from Robert Barkhouser and Steven Smee. They have
provided a report to the group, that report is scheduled for review on Jan
16. Second is a decision on CCD manufaturers. A recommendation has been
made to purcase an STA device rather than an E2V device. We anticipate
that both the optics design and CCD decisions will be made by the end of
January, prior to the next meeting. Any user input would be valuable
before then. Joanne requested to be included in the optics review. It
was decided to send the review information to all of the users committee
reps.
o ARCSAT update:
Bill gave a short report on the status of the telescope. The new user
interface software is working well, and the telescope upgrades are nearly
complete except for some remaining issues with dome control. Jim
Davenport and Leslie Hebb will be acting as beta testers providing
feedback to the new software and telescope performance. Suzanne said that
we are on target for beginning shared-risk scheduling in Quarter 2.
Institutions will not be charged for shared-risk time. Proposals will not
be as formal as on the 3.5-m. We anticipate a call for proposals at the
beginning of March. The February users meeting should have more
information on this.
o GIFS in Q2:
Mark and Suzanne reported on recent discussions with Goddard. It was
decided to freeze the instrument development at its current state and
finish the TUI software interface before Q2. Observers would have Q2 and
Q3 to try using the instrument with TUI and we will see how much usage
there is. The Goddard team has offered to provide a training run,
including information on data reduction, at APO during Q2 or Q3.
Interested users should mention this in their proposals and we will try to
schedule a block of GIFS runs. 2014 likely will be the year when the
future of GIFS at APO is decided based on user interest and feedback, so
usersare encouraged to propose and participate in the GIFS activity this
year and provide input to the users committee. One user asked about the
maturity of the reduction software and was directed to contact Carol Grady
at GSFC.
Goddard would like to make additional improvements to the instrument but
we have asked them to wait until we get feedback from users on the
instrument in a stable state and with TUI available. Michael asked about
the improvements, which include a photon counting detector and new readout
electronics.
o New users info, training:
Current information is at www.apo.nmsu.edu/arc35m, see "Training" link.
Alaina Bradley is working on a webpage for new users. When a draft is
available, Suzanne will pass it on to the new institutions for comments
before making it public.
ACTION ITEMS:
[open from previous meetings]:
===> ACTION: Users Committee members to check the Q1 schedule to confirm
that their observing programs conform to the on-site training policy [see
http://www.apo.nmsu.edu/arc35m/TrainingInfo_ARC35m.html]
There was no feedback from members, so we assume all Q1 programs are
compliant. Users committee members and schedulers should continue to be
vigilant on this topic. This action is now closed.
===> ACTION: Users Committee members to think about the context and
themes for a 20th-anniversary celebration for the 3.5-m, to be discussed
at the next Users Committee telecon.
The dates of the celebration will likely be May 12-13, 2014. An evening
reception and telescope viewing will take place on May 12, and a full day
of historical and science talks, along with thanks to past and present
contributors to the success of the telescope is planned on May 13. An
invitation list is being prepared, any users that would like to attend
should contact their users committee representative. This action will
remain open for discussion at the February meeting.
===> ACTION: Find users to assist with instrument documentation.
Status: We received some comments and have implemented them.
Action closed.
===> ACTION: Users committee members to query their communities for
interest in and volunteers to help with setting up a 3.5m user wiki.
Status: Jon Holtzman is working on this. Action closed.
[new from this meeting]:
===> ACTION ITEM: Committee members please query their users about issues
with reducing data due to the new format of the timestamp.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 10
February at 8:30 AM Pacific Time. The agenda and other materials will be
sent to the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1325 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO